# Manipulating Spectral Windings and Skin Modes through Nonconservative Couplings

### v1

@@ Abstract
The discovery of the non-Hermitian skin effect (NHSE) has revolutionized our understanding of wave propagation in non-Hermitian systems, highlighting unexpected localization effects beyond conventional theories. Here, we discover that NHSE, accompanied by multi-type spectral phases, can be induced by manipulating nonconservative couplings. By characterizing the spectrum through the windings of the energy bands, we demonstrate that band structures with identical, opposite, and even twisted windings can be achieved. These inequivalent types of spectra originate from the multi-channel interference resulting from the interplay between conservative and nonconservative couplings. Associated with the multi-type spectra, unipolar and bipolar NHSE with different eigenmode localizations can be observed. Additionally, our findings link the nonreciprocal transmission properties of the system to multiple spectral phases, indicating a connection with the skin modes. This work paves new pathways for investigating non-Hermitian topological effects and manipulating nonreciprocal energy flow.
Introduction.— Recent advancements in the study of non-Hermitian systems have garnered extensive interest due to their unique attributes and potential applications that extend beyond the realm of traditional Hermitian physics [1, 2, 3, 4, 5]. Non-Hermitian Hamiltonians, typically employed to describe systems that exchange energy with their environment, have revealed a number of intriguing phenomena [6, 7, 8, 9, 10]. Among these, the non-Hermitian skin effect (NHSE) results in localization of the eigenstates at the edges of the system, challenging conventional bulk-boundary correspondence and having profound theoretical and experimental implications [11, 12, 13, 14, 15, 16, 17]. In fact, characteristic signatures of the NHSE are directly reflected in the system’s spectrum, which displays complex band structures forming closed loops with nonzero windings in the quasi-energy space under periodic boundary conditions (PBC), distinctly different from those observed under open boundary conditions (OBC) [18, 19]. Experimentally, the NHSE has been demonstrated in various practical setups including optical systems [20, 21], acoustic systems [22, 23], and topolectrical circuits [24, 25]. These studies have potential applications in directional amplifiers [26, 27], enhanced sensors [28, 29], and efficient energy harvesting [30], showcasing the real-world utility of this effect.
@@
Traditionally, the realization of bipolar NHSE necessitates asymmetric coupling strengths that extend beyond nearest-neighbor couplings [38, 39, 40]. Here, we propose a different mechanism to realize and manipulate spectral windings and skin modes, namely through nonconservative couplings. Different from coherent couplings, nonconservative couplings connect elements of the systems indirectly through intermediates with gain or loss, exhibiting non-Hermitian characteristics in the phase terms of the couplings [41, 42, 43, 44, 45]. By considering a chain of resonance modes with conservative nearest-neighbor couplings and nonconservative next-nearest-neighbor couplings, we find that diverse spectra can appear by manipulating the nonconservative coupling phases. These include closed loops with the same and opposite windings, as well as twisted windings with two oppositely oriented loops, which arise by combining the nonconservative couplings with multi-channel interference provided by the chain. Associated with the multi-type spectra, unipolar and bipolar NHSE with different eigenmode localizations can be observed. To connect with the nonreciprocity performance, we also present evidence for a close link between the nonreciprocal transmission of the chain and multi-type spectral phases as well as NHSE.
Figure 1: (a) Schematic of the non-Hermitian chain with conservative nearest-neighbor couplings (solid lines) and nonconservative next-nearest-neighbor couplings ($ve^{i\phi_{a,b}}$, dashed lines). (b-c) Possible combinations of the upper and lower triangular plaquettes with the nonreciprocal flow. The direction of the nonreciprocal flow in each triangular plaquette (illustrated by the red and blue arrows) can be controlled by tuning the sign of the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$.
Model.— We consider a chain of bosonic modes with conservative nearest-neighbor couplings and nonconservative next-nearest-neighbor couplings. The system Hamiltonian is given by
@@
where $2\delta$ is the difference of the mode resonance frequencies, $w$ and $u$ denote the conservative nearest-neighbor intra- and inter-hopping terms, respectively. The non-Hermiticity of the system is induced by considering nonzero phases $\phi_{a}$ and $\phi_{b}$ of the nonconservative next-nearest-neighbor couplings, where the energy flow in the coupling process $\hat{a}_{j}^{\dagger}\hat{a}_{j+1}+\hat{a}_{j+1}^{\dagger}\hat{a}_{j}$ ($\hat{b}_{j}^{\dagger}\hat{b}_{j+1}+\hat{b}_{j+1}^{\dagger}\hat{b}_{j}$) are not conserved (an example to implement the nonconservative couplings can be found in the Supplemental Material [46]). As illustrated in Fig. 1(a), this periodic coupling structure can be interpreted as a combination of the upper ($a_{i}\leftrightarrow b_{i}\leftrightarrow a_{i+1}$) and lower ($b_{i}\leftrightarrow a_{i+1}\leftrightarrow b_{i+1}$) triangular plaquettes. In each triangular plaquette, nonreciprocal energy flow can be realized by combining nonconservative couplings with two-channel interference provided by the nonzero effective flux [45]. The direction of the nonreciprocal flow in the triangular plaquette can be changed by tuning the sign of the nonconservative coupling phase $\phi_{a}$ ($\phi_{b}$). When combining the upper and lower triangular plaquettes, the nonreciprocal flow can be both clockwise (anti-clockwise) by choosing $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})$, or in the opposite direction when $\text{sgn}(\phi_{a})=-\text{sgn}(\phi_{b})$, as shown in Fig. 1(b) and (c).
Multi-type spectra and NHSE.— Under the PBC, the Hamiltonian in Eq. (1) can be diagonalized in the momentum basis with $\hat{h}(k)=\langle k|\hat{H}|k\rangle=\sum_{a_{j},b_{j}}\langle k,a_{j}|\hat{H}|k,b_{j}\rangle\cdot|a_{j}\rangle\langle b_{j}|$, where $\hat{h}(k)$ denotes the momentum Hamiltonian. As shown in Fig. 2, the energy spectra under the PBC form loops in the complex quasi-energy space, which is distinct from the two open arcs under the OBC, indicating the emergence of the NHSE in our system [47]. To further investigate the energy spectra and wavefunctions of our model, we characterize the spectral phases by describing the band windings [18]. For any base point $E_{b}$, the winding number of the energy spectrum is given by
$W=\frac{1}{2\pi i}\int_{-\pi}^{\pi}\partial_{k}\text{ln det}[\hat{h}(k)-E_{b}\cdot\mathbb{I}_{2\times 2}]dk,$ (2)
@@
Figure 2: (a)-(b) Phase diagram of the spectral winding number for the phase transition of two energy bands, respectively. The red region indicates the winding number $W=1$ and the blue region indicates $W=-1$. The light yellow area represents the appearance of twisted spectral winding. (c)-(f) Spectra under PBC (red and blue for mode $a$ and $b$) and OBC (black points) of the model, with different non-conservative coupling phases $\phi_{b}=\pi/2,\,-\pi/2,\,-\pi/8$ and $\phi_{a}=\pi/2$ (c)-(e), and $\phi_{b}=-19\pi/20$, $\phi_{a}=\pi/20$ (f), respectively. Clockwise and counterclockwise winding directions of the PBC spectrum versus the quasi-momentum $k$, as indicated by the arrows, correspond to OBC skin modes localized on the left and right, respectively. Other parameters are $\delta=2,\,w=1,\,u=i,\,v=1$.
Characterizing the spectral topology by winding number $W$, we observe that the spectral phase diagrams of the two energy bands exhibit distinct behaviors when the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$ are adjusted, as depicted in Fig. 2 (a) and (b). The red and blue regions represent the closed-loop-like band structure with winding numbers $W=1$ and $W=-1$ [Fig. 2(c) and (d)], respectively. The light yellow area indicates the presence of the twisted winding [Fig. 2(e) and (f)]. To elucidate the characteristics of spatial profiles corresponding to different spectral phases, we analyze the spatial profiles of all eigenmodes under the OBC, illustrated in Fig. 3. Based on the eigenmode distributions across the lattice sites in our model, these eigenmodes are categorized into two groups: those predominantly localized on sites $a_{j}$ with minimal occupation on $b_{j}$, and vice versa. In the cases where $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})$, the two bands typically exhibit opposite circling directions, as depicted in Fig. 2. Under these conditions, the eigenmodes predominantly occupying sites $a_{j}$ localize toward the left edge, while the eigenmodes concentrated on sites $b_{j}$ localize at the right boundary, simultaneously. This arrangement is indicative of the bipolar NHSE, as illustrated in Fig. 3 (a) and (b). Conversely, when $\text{sgn}(\phi_{a})=-\text{sgn}(\phi_{b})$, the two bands share the same winding number and circle in the same direction. This scenario results in all eigenmodes localizing at the same boundary, characteristic of the conventional unipolar NHSE, as shown in Fig. 3 (c)and (d).
@@
Figure 3: (a)-(h) Distributions of all eigenmodes corresponding to the phase ①-④ in Fig. 2 respectively. Blue (red) lines represent the distributions of the eigenstates whose eigenvalue localized in the blue (red) loop of the PBC spectral on the $b$ ($a$) sites. Insets show the distributions of eigenmodes on the $a$ ($b$) sites, indicating a tiny population on the opposite side. Other parameters are $\delta=2,\,w=1,\,u=i,\,v=1,\,N=60$.
@@
The multi-type spectra and the NHSE localization transitions can also be observed in the context of the nonreciprocal flow in the periodic structure shown in Fig. 1. The band twisting is associated with the effective couplings between the upper and lower triangular plaquettes, specifically the energy flow between modes $a_{j+1}$ and $b_{j}$. When the signs of $\phi_{a}$ and $\phi_{b}$ are the same, the opposite nonreciprocal flow from modes $a_{j}$ to $a_{j+1}$ and from $b_{j+1}$ to $b_{j}$ corresponds to the presence of the bipolar skin effects for $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})$. In this scenario, the two nonreciprocal transmission channels between modes $a_{j+1}$ and $b_{j}$, provided by the upper and lower triangular plaquettes, interfere destructively. This results in negligible energy flow between modes $a_{j+1}$ and $b_{j}$, keeping the two bands relatively independent. Only when there is a significant difference between the phases $\phi_{a}$ and $\phi_{b}$, which results in a non-negligible effective coupling between the upper and lower triangular plaquettes, a band twist emerges. Conversely, when the signs of $\phi_{a}$ and $\phi_{b}$ differ, the nonreciprocal flows align, causing the same transmission from modes $b_{j}$ to $a_{j+1}$ in both the upper and lower triangular plaquettes. Constructive interference between these two transmission channels enhances the energy flow between the modes $b_{j}$ and $a_{j+1}$, facilitating the twisting of the bands with even a small difference in the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$.

### v2

@@ Abstract
The discovery of the non-Hermitian skin effect (NHSE) has revolutionized our understanding of wave propagation in non-Hermitian systems, highlighting unexpected localization effects beyond conventional theories. Here, we discover that NHSE, accompanied by multi-type spectral phases, can be induced by manipulating nonconservative couplings. By characterizing the spectra through the windings of the energy bands, we demonstrate that band structures with identical, opposite, and even twisted windings can be achieved. These inequivalent types of spectra originate from the multi-channel interference resulting from the interplay between conservative and nonconservative couplings. Associated with the multi-type spectra, unipolar and bipolar NHSE with different eigenmode localizations can be observed. Additionally, our findings link the nonreciprocal transmission properties of the system to multiple spectral phases, indicating a connection with the skin modes. This paper paves new pathways for investigating non-Hermitian topological effects and manipulating nonreciprocal energy flow.
I Introduction
Recent advancements in the study of non-Hermitian systems have garnered extensive interest due to their unique attributes and potential applications that extend beyond the domain of traditional Hermitian physics [1, 2, 3, 4, 5]. Non-Hermitian Hamiltonians, typically employed to describe systems that exchange energy with their environment, have revealed a number of intriguing phenomena [6, 7, 8, 9, 10, 11, 12]. Among these, the non-Hermitian skin effect (NHSE) results in localization of the eigenstates at the edges of the system, challenging conventional bulk-boundary correspondence and having profound theoretical and experimental implications [13, 14, 15, 16, 17, 18, 19]. In fact, characteristic signatures of the NHSE are directly reflected in the system’s spectrum, which displays complex band structures forming closed loops with nonzero windings in the quasi-energy space under periodic boundary conditions (PBC), distinctly different from those observed under open boundary conditions (OBC) [20, 21]. Experimentally, the NHSE has been demonstrated in various practical setups including optical systems [22, 23], acoustic systems [24, 25], and topolectrical circuits [26, 27]. These studies have potential applications in directional amplifiers [28, 29], enhanced sensors [30, 31], and efficient energy harvesting [32], showcasing the real-world utility of this effect.
@@ I Introduction
Traditionally, the realization of bipolar NHSE necessitates asymmetric coupling strengths that extend beyond nearest-neighbor couplings [40, 41, 42, 43]. Here, we propose a different mechanism to realize and manipulate spectral windings and skin modes, namely through nonconservative couplings. Different from coherent couplings, nonconservative couplings connect elements of the systems indirectly through intermediates with gain or loss, exhibiting non-Hermitian characteristics in the phase terms of the couplings [44, 45, 46, 47, 48]. By considering a chain of resonance modes with conservative nearest-neighbor couplings and nonconservative next-nearest-neighbor couplings, we find that diverse spectra can appear by manipulating the nonconservative coupling phases. These include closed loops with the same and opposite windings, as well as twisted windings with two oppositely oriented loops, which arise by combining the nonconservative couplings with multi-channel interference provided by the chain. Associated with the multi-type spectra, unipolar and bipolar NHSE with different eigenmode localizations can be observed. To connect with the nonreciprocity performance, we also present evidence for a close link between the nonreciprocal transmission of the chain and multi-type spectral phases as well as NHSE.
The rest of this paper is organized as follows. In Sec. II, we introduce the system model of a non-Hermitian chain with nonconservative couplings. In Sec. III, we present the spectral phase diagram, eigenstates distributions and the generalized Brillouin zone (GBZ) to illustrate the multi-type spectra and NHSE. We also show the influence of the phase and strength of the nonconservative couplings on the spectral twisting phenomenon. In Sec. IV, we present evidence for a close relationship between the multi-type spectral phases and the nonreciprocal transmission of the chain. Finally, we summarize our results in Sec. V. In the Appendixes, we provide several parts of the detailed derivations, including the derivation of the effective Hamiltonian with nonconservative couplings (Appendix A), the phase diagram of two-band winding number and Inverse Participation Ratio (IPR) of the model (Appendix B) and the nonreciprocal energy transmission (Appendix C).
Figure 1: (a) Schematic of the non-Hermitian chain with conservative nearest-neighbor couplings (solid lines) and nonconservative next-nearest-neighbor couplings ($ve^{i\phi_{a,b}}$, dashed lines). Red balls represent subchain $a_{j}$ while blue balls represent subchain $b_{j}$. (b-c) Possible combinations of the upper and lower triangular plaquettes with the nonreciprocal flow. The direction of the nonreciprocal flow in each triangular plaquette (illustrated by the red and blue arrows) can be controlled by tuning the sign of the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$.
II Model
We consider a chain of bosonic modes with conservative nearest-neighbor couplings and nonconservative next-nearest-neighbor couplings. The system Hamiltonian is given by
@@ II Model
where $2\delta$ is the difference of the mode resonance frequencies, and $w$ and $u$ denote the conservative nearest-neighbor intra- and inter-hopping terms, respectively. The non-Hermiticity of the system is induced by considering nonzero phases $\phi_{a}$ and $\phi_{b}$ of the nonconservative next-nearest-neighbor couplings, where the energy flow in the coupling process $\hat{a}_{j}^{\dagger}\hat{a}_{j+1}+\hat{a}_{j+1}^{\dagger}\hat{a}_{j}$ ($\hat{b}_{j}^{\dagger}\hat{b}_{j+1}+\hat{b}_{j+1}^{\dagger}\hat{b}_{j}$) is not conserved (an example to implement the nonconservative couplings can be found in the Appendix A). As illustrated in Fig. 1(a), this periodic coupling structure can be interpreted as a combination of the upper ($a_{i}\leftrightarrow b_{i}\leftrightarrow a_{i+1}$) and lower ($b_{i}\leftrightarrow a_{i+1}\leftrightarrow b_{i+1}$) triangular plaquettes. In each triangular plaquette, nonreciprocal energy flow can be realized by combining nonconservative couplings with two-channel interference provided by the nonzero effective flux [48]. The direction of the nonreciprocal flow in the triangular plaquette can be changed by tuning the sign of the nonconservative coupling phase $\phi_{a}$ ($\phi_{b}$). When combining the upper and lower triangular plaquettes, the nonreciprocal flow can be the same direction by choosing $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})$, or in the opposite direction when $\text{sgn}(\phi_{a})=-\text{sgn}(\phi_{b})$, as shown in Fig. 1(b) and (c).
III Multi-type spectra and NHSE
III.1 Spectral phase diagram and distribution of eigenstates
Under the PBC, the Hamiltonian in Eq. (1) can be diagonalized in the momentum basis with $\hat{h}(k)=\langle k|\hat{H}|k\rangle=\sum_{a_{j},b_{j}}\langle k,a_{j}|\hat{H}|k,b_{j}\rangle|a_{j}\rangle\langle b_{j}|$, where $\hat{h}(k)$ denotes the momentum Hamiltonian. As shown in Fig. 2, the energy spectra under the PBC form loops in the complex quasi-energy space, which is distinct from the two open arcs under the OBC, indicating the emergence of the NHSE in our system [49]. To further investigate the energy spectra and wavefunctions of our model, we characterize the spectral phases by describing the band windings [20]. For any base point $E_{b}$, the winding number of the energy spectrum is given by
$W=\frac{1}{2\pi i}\int_{-\pi}^{\pi}\partial_{k}\text{ln det}[\hat{h}(k)-E_{b}\mathbb{I}_{2\times 2}]dk,$ (2)
@@ III.1 Spectral phase diagram and distribution of eigenstates
Figure 2: (a)-(b) Phase diagram of the spectral winding number for the phase transition of two energy bands, respectively. The red region indicates the winding number $W=1$ and the blue region indicates $W=-1$. The light yellow area represents the appearance of twisted spectral winding. (c)-(f) Spectra under PBC (red and blue for mode $a$ and $b$) and OBC (black points) of the model, with different non-conservative coupling phases $\phi_{b}=\pi/2,\,-\pi/2,\,-\pi/8$ and $\phi_{a}=\pi/2$ (c)-(e), and $\phi_{b}=-19\pi/20$, and $\phi_{a}=\pi/20$ (f), respectively. Clockwise and counterclockwise winding directions of the PBC spectrum vs the quasi-momentum $k$, as indicated by the arrows, correspond to OBC skin modes localized on the left and right, respectively. The system size is $N=100$ and other parameters are $w/\delta=1/2,\,u/\delta=i/2,\,v/\delta=1/2$.
Characterizing the spectral topology by winding number $W$, we observe that the spectral phase diagrams of the two energy bands exhibit distinct behaviors when the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$ are adjusted, as depicted in Fig. 2 (a) and (b). The red and blue regions represent the closed-loop-like band structure with winding numbers $W=1$ and $-1$ [Fig. 2(c) and (d)], respectively. The light yellow area indicates the presence of the twisted winding [Fig. 2(e) and (f)]. To elucidate the characteristics of spatial profiles corresponding to different spectral phases, we analyze the spatial profiles of all eigenmodes under the OBC, illustrated in Fig. 3. Based on the eigenmode distributions across the lattice sites in our model, these eigenmodes are categorized into two groups: those predominantly localized on sites $a_{j}$ with minimal occupation on $b_{j}$, and vice versa. In the cases where $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})$, the two bands typically exhibit opposite circling directions, as depicted in Fig. 2. Under these conditions, the eigenmodes predominantly occupying sites $a_{j}$ localize toward the left edge, while the eigenmodes concentrated on sites $b_{j}$ localize at the right boundary, simultaneously. This arrangement is indicative of the NHSE, as illustrated in Fig. 3 (a) and (b). Conversely, when $\text{sgn}(\phi_{a})=-\text{sgn}(\phi_{b})$, the two bands share the same winding number and circle in the same direction. This scenario results in all eigenmodes localizing at the same boundary, characteristic of the conventional unipolar NHSE, as shown in Fig. 3 (c)and (d).
III.2 Twisted winding
@@
Figure 3: (a)-(h) Distributions of all eigenmodes corresponding to phase 1-4 in Fig. 2 respectively. Blue (red) lines represent the distributions of the eigenstates with eigenvalue localized in the blue (red) loop of the PBC spectrum on the $b$ ($a$) sites. Insets show the distributions of eigenmodes on the $a$ ($b$) sites, indicating a tiny population on the opposite side. The system size is $N=100$ and other parameters are $w/\delta=1/2,\,u/\delta=i/2,\,v/\delta=1/2$.
@@
The multi-type spectra and the NHSE localization transitions can also be observed in the context of the nonreciprocal flow in the periodic structure shown in Fig. 1. The closed-loop-like band structure with a nonzero winding number indicates the unipolar localization skin effects consistent with the unidirectional reciprocal transmission. However, twisted winding represents the respectively bidirectional localization skin effects in the corresponding subchain $a_{j}$ or $b_{j}$, which indicates competing energy transport in different directions. In our model, the main competition occurs between the inter-chain and intra-chain energy transmission channels, which may have different directions. Specifically, taking chain $a_{j}$ (red balls) as an example, the intra-chain energy transfer direction among modes $a_{j}$ may differ from the inter-chain transfer direction between $a_{j}$ and $b_{j}$ (or $a_{j+1}$ and $b_{j}$). When these two transfer directions are different and the inter-chain energy transfer is non-negligible, it will affect the localization direction of mode $a_{j}$, corresponding to the emergence of the twisted winding. When the signs of $\phi_{a}$ and $\phi_{b}$ are the same, the opposite nonreciprocal flow from modes $a_{j}$ to $a_{j+1}$ and from $b_{j+1}$ to $b_{j}$ corresponds to the presence of the bipolar skin effects for $\text{sgn}(\phi_{a})=\text{sgn}(\phi_{b})$. In this scenario, the two inter-chain transmission channels between modes $a_{j+1}$ and $b_{j}$, provided by the upper and lower triangular plaquettes, interfere destructively. This results in negligible energy flow between modes $a_{j+1}$ and $b_{j}$, keeping the two bands relatively independent. Only when there is a significant difference between the phases $\phi_{a}$ and $\phi_{b}$, which results in a non-negligible effective coupling between the upper and lower triangular plaquettes, a band twist emerges. Conversely, when the signs of $\phi_{a}$ and $\phi_{b}$ differ, the nonreciprocal flows differently, causing the same transmission from modes $b_{j}$ to $a_{j+1}$ in both the upper and lower triangular plaquettes. Constructive interference between the two inter-chain transmission channels enhances the energy flow between the modes $b_{j}$ and $a_{j+1}$, facilitating the twisting of the bands with even a small difference in the nonconservative coupling phases $\phi_{a}$ and $\phi_{b}$.
Moreover, we investigate the influence of the nonconservative coupling strength on the phase diagram of the spectral winding number. In Fig. 4, we present the phase diagram of the spectral winding number for the phase transition of a single energy band. The red region indicates the winging number $W=1$ and the blue region indicates $W=-1$. The light yellow area represents the appearance of twisted spectral winding. As the nonconservative coupling strength $v/\delta$ increases from 0.1 to 0.4, we observe an expansion of the twisting region in the phase diagram. This expansion can be attributed to the increasing competition between the inter-chain and intra-chain energy transmission in the system.
Figure 4: (a)-(d) Phase diagram of the spectral winding number for the phase transition of a single energy band with $w/\delta=1/2,\,u/\delta=i/2$. The red region indicates the winding number $W=1$ and the blue region indicates $W=-1$. The light yellow area represents the appearance of twisted spectral winding. The strength of the nonconservative coupling $v/\delta$ increases from 0.1 to 0.4.
III.3 GBZ
[…]
The properties of the non-Hermitian skin effect can also be characterized by the GBZ [13], which is calculated by
$f(\beta,E)=\text{det}[E\cdot\mathbb{I}_{2\times 2}-h(\beta)]=0,$ (3)
with $h(\beta)=h(e^{ik}\rightarrow\beta)$ and $E$ being the eigenvalues under the OBC. Here, $f(\beta,E)$ is a quartic equation for $\beta$, having four solutions $\beta_{i}$, with $|\beta_{1}|\leqslant|\beta_{2}|\leqslant|\beta_{3}|\leqslant|\beta_{4}|$. The trajectory of solution $\beta$ satisfying $|\beta_{2}|=|\beta_{3}|$ gives rise to the GBZ.
Fig. 5 shows the GBZs for the different nonconservative coupling phases. When the nonconservative coupling phases $\phi_{b}\ =\ \pi/2$ and $\phi_{a}=\pi/2$, the spectra winding numbers of the two bands are $W_{1}=1$ and $W_{2}=-1$. As seen in Fig. 5(a), one set of sub-GBZs with $\text{Re}(E)>0$ is located inside the unit circle (blue dashed lines), indicating localization on the left boundary. However, another set of sub-GBZs with $\text{Re}(E)<0$ lies outside the unit circle, indicating localization on the right side. When $\phi_{b}\ =\ -\pi/2$ and $\phi_{a}=\pi/2$, all spectra winding numbers are $W_{1}=W_{2}=1$, indicating that all skin modes are localized at the left boundary. The corresponding GBZs are located inside the unit circle, as shown in Fig. 5(b). In the twisted winding region [Fig. 5 (c) and (d)], GBZs intersect with the unit circle. These intersection points are Bloch points, indicating that the corresponding eigenstates of the system exhibit Bloch-wave-like behavior.
IV Connection with nonreciprocity